Suppose two rats A and B have been trained to navigate a large maze.
X=Time taken by rat A to navigate the maze X~N(80,100)
Y=Time taken by rat B to navigate the maze, Y~N(78,169)
On any given day, what is the probability that rat A runs the maze faster than rat B?
1
Expert's answer
2019-09-30T09:35:25-0400

We will suppose that "X" and "Y" are independent variables. So we need to find"P(X<Y)=\iint\limits_{x<y}\rho_X(x)\rho_Y(y)dxdy=" "\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^y\rho_X(x)\rho_Y(y)dxdy=" "=\frac{1}{100\sqrt{2\pi}}\frac{1}{169\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^ye^{-\frac{(x-80)^2}{2*100^2}}e^{-\frac{(y-78)^2}{2*169^2}}dxdy".Let "u=\frac{x-80}{100}", then "\frac{1}{100\sqrt{2\pi}}\frac{1}{169\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^ye^{-\frac{(x-80)^2}{2*100^2}}e^{-\frac{(y-78)^2}{2*169^2}}dxdy=" "=\frac{1}{100\sqrt{2\pi}}\frac{1}{169\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{\frac{y-80}{100}}e^{-\frac{u^2}{2}}e^{-\frac{(y-78)^2}{2*169^2}}d(100u+80)dy=" "=\frac{1}{\sqrt{2\pi}}\frac{1}{169\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{\frac{y-80}{100}}e^{-\frac{u^2}{2}}e^{-\frac{(y-78)^2}{2*169^2}}dudy".Let "v=\frac{y-78}{169}" , then "\frac{y-80}{100}=\frac{169v-2}{100}" and "\frac{1}{\sqrt{2\pi}}\frac{1}{169\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{\frac{y-80}{100}}e^{-\frac{u^2}{2}}e^{-\frac{(y-78)^2}{2*169^2}}dudy=" "=\frac{1}{\sqrt{2\pi}}\frac{1}{169\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{\frac{169v-2}{100}}e^{-\frac{u^2}{2}}e^{-\frac{v^2}{2}}dud(169v+78)=" "=\frac{1}{\sqrt{2\pi}}\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{\frac{169v-2}{100}}e^{-\frac{u^2}{2}}e^{-\frac{v^2}{2}}dudv"


We have the following region of integrate in image. Make orthogonal change of variables, where the line "u=\frac{169v}{100}" will be the axis "u'" and orthogonal line "u=-\frac{100}{169}v" will be axis "v'" .

So "\begin{pmatrix}
 v \\
 u
\end{pmatrix}=
\begin{pmatrix}
 \cos\alpha & -\sin\alpha \\
 \sin\alpha & \cos\alpha
\end{pmatrix}
\begin{pmatrix}
 v' \\
 u'
\end{pmatrix}" , where "\tan\alpha=-\frac{100}{169}" , that is "\cos\alpha=\frac{169}{\sqrt{38561}}" and "\sin\alpha=-\frac{100}{\sqrt{38561}}" .

We obtain that line "u=\frac{169v-2}{100}" will be transformed into "-\frac{100}{\sqrt{38561}}v'+\frac{169}{\sqrt{38561}}u'=\frac{169}{100}\bigl(\frac{169}{\sqrt{38561}}v'+\frac{100}{\sqrt{38561}}u'\bigr)-\frac{1}{50}" or "v'=\frac{2}{\sqrt{38561}}"

Line "u=-\frac{100}{169}v" will be transformed into "u'=0".

We obtain the new region of integration "\bigl\{(u',v'): v'\ge\frac{2}{\sqrt{38561}}\bigr\}"

Jacobian of our transformation is "J=\bigl|\begin{pmatrix}
 \cos\alpha & -\sin\alpha \\
 \sin\alpha & \cos\alpha
\end{pmatrix}\bigr|=1"

Since our transformation is orthogonal, we have "u^2+v^2=(u')^2+(v')^2" , so "\frac{1}{\sqrt{2\pi}}\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty}\int\limits_{-\infty}^{\frac{169v-2}{100}}e^{-\frac{u^2}{2}}e^{-\frac{v^2}{2}}dudv="

"=\frac{1}{\sqrt{2\pi}}\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty}\int\limits_{\frac{2}{\sqrt{38561}}}^{+\infty}Je^{-\frac{(u')^2}{2}}e^{-\frac{(v')^2}{2}}du'dv'=" "=\frac{1}{\sqrt{2\pi}}\int\limits_{\frac{2}{\sqrt{38561}}}^{+\infty}e^{-\frac{(u')^2}{2}}du'\cdot\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty}e^{-\frac{(v')^2}{2}}dv'=" "=\Bigl(1-\Phi\bigl(\frac{2}{\sqrt{38561}}\bigr)\Bigr)\Phi(+\infty)=1-\Phi\bigl(\frac{2}{\sqrt{38561}}\bigr)" , where "\Phi(x)=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^x e^{-\frac{y^2}{2}}dy"

https://www.math.arizona.edu/~rsims/ma464/standardnormaltable.pdf - Normal Disribution Table

Since "\frac{2}{\sqrt{38561}}\approx 0.01" , from table we see that "\Phi\bigl(\frac{2}{\sqrt{38561}}\bigr)\approx\Phi(0.01)\approx 0.50399" , so "1-\Phi\bigl(\frac{2}{\sqrt{38561}}\bigr)\approx 0.49601"

Answer: "0.49601"